OpenStudy (anonymous):

Calculate the density of a solid substance if a cube measuring 2.72 cm on one side has a mass of 130g.

OpenStudy (unklerhaukus):

density \(\varrho\), is equal to the mass \(m\), per volume \(V\) \[\varrho = m/V\] Calculate the volume of the cube

OpenStudy (anonymous):

How do you calculate the volume of the cube?

OpenStudy (unklerhaukus):

the volume \(V\), of a cube is equal to the length of its sides \(s\), cubed. \[V =s^3\]

OpenStudy (anonymous):

so you would cube the 2.72?

OpenStudy (unklerhaukus):

you are given that the sides are \(s=2.72[\text{cm}]\)

OpenStudy (unklerhaukus):

What volume does this make?

OpenStudy (anonymous):

If you cube the 2.72 cm?

OpenStudy (unklerhaukus):

yes,

OpenStudy (anonymous):

20.123648 cm^3

OpenStudy (unklerhaukus):

good, thats the volume, now the density is \[\varrho = m/V = \frac{130[\text g]}{20.123648 [\text{cm}^3]}=\]

OpenStudy (anonymous):

6.460061317g/cm^3

OpenStudy (unklerhaukus):

yeah thats it, now you might want to round to 3 significant figures, (because the values given are themselves only accurate to 3sig.figs)

OpenStudy (anonymous):

So 6.460

OpenStudy (unklerhaukus):

(that is four significant figures)

OpenStudy (anonymous):

Whoops 6.46

OpenStudy (unklerhaukus):

yeah, i would leave my final answer as \(\varrho=6.46\,[\text g/\text{cm}^3]\)

OpenStudy (anonymous):

Alright thank you. There is another question like it wanting to calculate the mass of a cube of the same substance measuring 7.51cm on one side.

OpenStudy (unklerhaukus):

OK, well since the substance is the same, the density (an intrinsic quality) will remain the same

OpenStudy (unklerhaukus):

rearranging\[\varrho=m/V\] we get\[m =\varrho\cdot V\] again the volume of a cube is \(s^3\)

OpenStudy (anonymous):

7.51^3 = 423.564751

OpenStudy (unklerhaukus):

now multiply that, by the density we found earlier

OpenStudy (anonymous):

2736.228291

OpenStudy (unklerhaukus):

and what are the units?

OpenStudy (anonymous):

grams

OpenStudy (anonymous):

since mass was being calculated for

OpenStudy (unklerhaukus):

yeah, so what is our final answer with units (and to 3sig.figs)

OpenStudy (anonymous):

Alright and if it wants it in scientific notation would it be 2.73 x 10^3?

OpenStudy (unklerhaukus):

\[2.74 \times 10^3\,[\text g]=2.74\,[\text{kg}]\]

OpenStudy (anonymous):

Thanks. Didn't notice the 6 there.

OpenStudy (anonymous):

Thanks for the help!

OpenStudy (unklerhaukus):

notice that the ratio of volumes \(424\,[\text{cm}^3]/20.1\,[\text{cm}^3]\approx21\), is equal to the ratio of masses \(2740\,[\text g]/130\,[\text g]\approx21\)

OpenStudy (unklerhaukus):

(as we should expect for a substance of constant density )

OpenStudy (anonymous):

I didn't notice that. That's interesting and that does make sense for it to.